1) Match these statements with the correct terms below them. Each statement will only be used once!

incident wave

wave front

medium

wave

wave train

reflected wave


  1. A disturbance that moves outwards from its point of origin, transferring energy through the medium
  2. An imaginary line that joins all points reached by the wave at the same instant 
  3. The material through which the wave travels 
  4. A wave front moving from the point of origin toward a barrier
  5. A wave front moving away from a barrier 
  6. A series of waves forming a continuous series of crests and troughs


2) Sound waves travel through seawater at about 1500m/s. The frequency that would generate a wavelength of 1.55 m in seawater is ___ Hz. 


3) A speaker system generates sound waves at a frequency of 2200 Hz. If the wave speed in air is 325 m/s, the wavelength is ___m



1
Expert's answer
2021-07-20T09:46:28-0400

"\text{Answer:}"

"\text{1. wave}\newline
\text{A disturbance that moves outwards from its point of origin, }\newline
\text{transferring energy through the medium}"

"\text{2. wave front}"

"\text{An imaginary line that joins all points reached by the wave}\newline
\text{at the same instant }"

"3.\text{ medium}\newline"

"\text{The material through which the wave travels }"

"\text{4. incident wave

}"

"\text{A wave front moving from the point of origin toward a barrier}"

"\text{5. reflected wave}"

"\text{A wave front moving away from a barrier }"

"\text{6.wave front}"

"\text{A series of waves forming a continuous series }"

"\text{of crests and troughs}"


"\text{2)}"

"\lambda = \frac{v}{f};"

"f = \frac{v}{\lambda}=\frac{1500}{1.55}=967.74Hz"

"\text{Answer:}"

"\text{Sound waves travel through seawater at about 1500m/s. }"

"\text{The frequency that would generate a wavelength }"

"\text{of 1.55 m in seawater is 966.74 Hz.}"

"\text{3)}"

"\lambda = \frac{v}{f}= \frac{325}{2200}\approx0.15m"

"\text{Answer:}"

"\text{A speaker system generates sound waves at a frequency of 2200 Hz. }"

"\text{If the wave speed in air is 325 m/s,}"

"\text{the wavelength is 0.15m}"